OpenStudy (babybritt002):

Iron is sold in sheets that are 2.0 cm thick and 120.0 cm wide. You want to buy a sheet of iron with a mass of 50,000.0 g. What must be the length of the sheet of iron?

OpenStudy (anonymous):

I would say not enough information because you don't have the amount of grams for the given dimensions.

OpenStudy (aaronq):

you can solve this if you know the density of iron (which was likely in a previous problem for you). from google, d=7.874 g/cm^3 Now using the density formula, we know we need the volume and the mass. We were given a mass, but no volume, but two dimensions, we know the formula for volume for a rectangular prism is \(V=l*w*h\) (thickness is height here). so we sub in and solve for length. \(\sf d=\dfrac{m}{V}=\dfrac{m}{l*w*h}\rightarrow l=\dfrac{m}{d*w*h}=\dfrac{50,000.0 ~g}{7.874 g/cm^3*2 ~cm*120.0~cm }\)